OpenStudy (anonymous):

create and equation like this one (9y^5) / (3y^2) = (3^2 * y^5) / (3^1 * y^2) = 3^(2-1) * y^(5-2) = 3^1 * y^3 = 3y^3

OpenStudy (anonymous):

(9y^5) / (3y^2) = (3^2 * y^5) / (3^1 * y^2) = 3^(2-1) * y^(5-2) = 3^1 * y^3 = 3y^3

jhonyy9 (jhonyy9):

why you need to create an equation like this ?

OpenStudy (anonymous):

to get a example

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2}dx/(\tan (x))^{\sqrt{2}}\]